OpenStudy (anonymous):

If M(x, y) is the midpoint of segment AB with A(4, 6) and B(−8, −12), then which Quadrant is M located in? Quadrant I Quadrant II Quadrant III Quadrant IV .i got quadrant IV an i right?..

OpenStudy (anonymous):

iii

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

@Bailey, do u know how and why your ans is wrong???

OpenStudy (anonymous):

mid point formula says coordinates of mid point are x1 + x2 y1 + y2 ------- , --------- 2 2 in your case x1 + x2 = 4 + (-8) = -4 x1 + x2 -4 ------- = ---- = -2 2 2 y1 + y2 6 + (-12) -6 ------- = ---------- = ------ = -3 2 2 2 so your mid point coordinates are (-4, -3) i.e. both negatives so it will lie in the IIIrd quadrant